OpenStudy (anonymous):

1/6x + 3 = -6 what is x?

OpenStudy (anonymous):

give me a hug?

OpenStudy (luigi0210):

Welcome to Openstudy @Kristenkayy :)

OpenStudy (compassionate):

WELCOME TO OPENSTUDY! 1/6x + 3 = -6 1/6x = -9 Now, you can cancel x by switching to the opposite. Giving you -54.

OpenStudy (anonymous):

1. transpose 3 to the right side (or subtract both sides) 2. cross multiply ans simplify..

OpenStudy (luigi0210):

Nice job Comp

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\)

OpenStudy (luigi0210):

@DUDE..IM..A..DUCK stop freaking the new users out xD

OpenStudy (luigi0210):

I need to learn that latex thing..

OpenStudy (compassionate):

Indeed, you do.

OpenStudy (anonymous):

they want my love